OpenStudy (anonymous):

a cone with a radius of 8ft and a slant height of 17ft what is the surface area of this cone? Use 22\7 for pi.

OpenStudy (anonymous):

Facts: \[\text{Slant Height = 17}\] \[\pi = \frac{22}{7}\] \[r = 17\] \[\text{Surface area (not including the base) = }\pi*r*Slant Height\] \[\text{Area of the base = }\pi * r^{2}\]